OpenStudy (anonymous):

Use DeMoivre's Theorem: ((sqrt3)+(i))^3

OpenStudy (earthcitizen):

\[(r <\theta)^{n} = (r(\cos(\theta)+isin(\theta))^{n} = r ^{n}(\cos(n \theta) +isin(n \theta))\]

OpenStudy (earthcitizen):

\[\sqrt{3} +i = 2 \angle 30^{o}\]

OpenStudy (earthcitizen):

\[(r \angle \theta)^n = r^n \angle(n \theta) \]

OpenStudy (earthcitizen):

\[2 \angle 30^o =2^3 \angle (3*30^o) = 8 \angle 90^o \]

OpenStudy (anonymous):

8(cos(pi)+i sin(pi)) Is that right?

OpenStudy (earthcitizen):

it's in polar form

OpenStudy (earthcitizen):

pi/2

OpenStudy (anonymous):

Ahh...so confusing. sorry about this >_< So it would be 8(cos(pi/2)+i sin(pi/2))

OpenStudy (earthcitizen):

yup